OpenStudy (anonymous):

a = 16, b = 28, c = ?

Nnesha (nnesha):

right triangle ??

OpenStudy (anonymous):

If it is a right triangle, just do py thag!

OpenStudy (anonymous):

16^2+28^2=c^2

Nnesha (nnesha):

if yes then use Pythagorean theorem \[\huge\rm a^2 + b^2 = c^2\]  c is longest side of right triangle

OpenStudy (anonymous):

16^2+28^2= c^2 sqrt 1040= sqrt c^2 32.2=c

Nnesha (nnesha):

.

OpenStudy (anonymous):

what's the point of putting a period?? ^^

Nnesha (nnesha):

there is a reason of posting "."

OpenStudy (dtan5457):

@pamela.g Can't give direct answers, that's why

OpenStudy (anonymous):

Oh my bad.